ellaisyoung:

A set of numbers is shown below: {1.8, 3.8, 5.4, 8.4, 13.2} The number __ from the set will make the equation x + 4 = 7.8 true.

carmelle:

Subtract 4 from both sides 4 - 4 = ? 7.8 - 4 = ?

carmelle:

@ellaisyoung does that make sense or no-?

ellaisyoung:

umm no

carmelle:

ok, so we have to subtract 4 from both sides to isolate the variable(which is `x`).

carmelle:

do u understand now?

carmelle:

x + 4 = 7.8 - 4 - 4

ellaisyoung:

yeah so the answer is....

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ellaisyoung yeah so the answer is.... \(\color{#0cbb34}{\text{End of Quote}}\) we cant give direct answers, so now, subtract- `7.8 - 4`

ellaisyoung:

3.8?

carmelle:

yes, that means `x = 3.8`

carmelle:

The number `3.8` from the set will make the equation x + 4 = 7.8 true.

carmelle:

do you understand-?

ellaisyoung:

yeah thanks

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ellaisyoung yeah thanks \(\color{#0cbb34}{\text{End of Quote}}\) np